You are provided with the following demand and cost information for a monopolist.

P=60-0.25 Q and TC=Q2+40Q+50 Where P is price, Q is quantity of the product and TC is total cost. Determine the profit maximizing level of output and prices. (10Marks)


1
Expert's answer
2021-06-23T09:32:45-0400

"P = 60 -0.25Q"


"TR = P.Q"


"= (60-0.25Q)Q"


"TR = 60Q - 0.25Q^2"


"MR = \frac{dTR}{dQ}"


"= 60(1) - 0.5Q"


"MR = 60 - 0.5Q"


"TC = Q^2 + 40Q + 50"


"MC = \frac{dTC}{dQ}"


"= 2Q + 40"


Monopolistic profit-Maximizing condition


"MC = MR"


"60-0.5Q = 2Q + 40"


"60-40 = 2Q + 0.5Q"


"20 = 2.5Q"


"Q = \frac{20}{2.5 }"


"Q = 8"


a) Profit Maximizing price


Substituting into P equation


"P = 60-0.25Q"


"= 60 -0.25\times8"


"60-2"


Profit Maximizing price "= 58"


b) Profit Maximizing level of output


"P = MC"


"60-0.25(Q) = 2Q + 40"


"60-40 = 2Q +0.25Q"


"20 = 2.25Q"


"= \frac{20}{2.25}"


Profit Maximizing level of output "= 8.89"